# Algebraic K-theory of finite algebras over higher local fields

### v2

@@ 3. Negative topological cyclic homology
For a spectrum $A$, we define the height of $A$ to be the number $h\in\mathbb{N}$ such that for any $0\leq i\leq h$ one have $L_{K(i)}(A)\not\simeq 0,$ and $L_{K(h+1)}A\simeq 0$
@@
Since $\operatorname{BP}\langle n\rangle[x]/x^{e}$ is complex orientated, hence according to [DFHH14, Chapter 6, Proposition 4.1] the localization can be computed as $(v_{m}^{-1}\operatorname{BP}\langle n\rangle[x]/x^{e})^{\wedge}_{(p,v_{1},\ldots,v_{m-1})}$. Therefore for any $m\geq n+1$, $L_{K(m)}\operatorname{BP}\langle n\rangle[x]/x^{e}\simeq 0$ by the fact $v_{m}$ maps to 00 in homotopy, see Remark 2.27. According to Claim 2.9, there is an $\mathbb{E}_{3}$-algebra map $\operatorname{BP}\langle n\rangle[x]/x^{e}\to\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}$. Now by Claim 2.10, there is an $\mathbb{E}_{3}$-algebra map $\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}\to\operatorname{BP}\langle n\rangle$ induced by an $\mathbb{E}_{3}$-$\operatorname{MU}$-algebra map $\operatorname{MU}[x]/x^{e}\to\operatorname{MU}$, using Construction 2.3. Since $\operatorname{K}(-)$ is a lax symmetric monoidal functor [BGT13, Proposition 10.11], and $\operatorname{Mod}_{\operatorname{BP}\langle n\rangle[x]/x^{e}}\to\operatorname{Mod}_{\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}}\to\operatorname{Mod}_{\operatorname{BP}\langle n\rangle}$ are $\mathbb{E}_{2}$-monoidal maps between $\mathbb{E}_{2}$-monoidal categories [Lur17, Proposition 4.5.8.20]. We therefore have following $\mathbb{E}_{2}$-ring maps
@@
With the same notation as in Section 2, $\Lambda_{\operatorname{BP}\langle n\rangle}(x)$ is height $n$, $\operatorname{BP}\langle n\rangle/v_{n}^{e}$ is height $n-1$. And we have
@@
The second $\mathbb{E}_{1}$-$\operatorname{MU}^{hS^{1}}$-algebra map above is induced by an $\mathbb{E}_{1}$-$\operatorname{MU}$-algebra map and applying $-^{hS^{1}}$.
@@ 4. Segal conjecture
We investigate Lichtenbaum–Quillen phenomena in this section, we are reduce to prove Segal conjecture and weak canonical vanishing theorem for certain cyclotomic spectra due to Theorem 4.7. In particular, we prove Proposition 4.16, 4.20. First we review some basic concepts and related techniques.
@@
We say a spectrum $X$ is finite if $X$ is bounded ($\exists a,b\in\mathbb{Z},X\in\operatorname{Sp}_{[a,b]}$), and $\pi_{k}X$ is a finite group for any $k\in\mathbb{Z}$. A spectrum $X$ is called fp-spectrum if $X$ is bounded below, $p$-complete, and there exists a finite complex $F$ such that $F\otimes X$ is finite in the previous sense.
@@
The ring spectra $(\operatorname{BP}\langle n\rangle[x]/x^{e})^{\wedge}_{p},(\Lambda_{\operatorname{BP}\langle n\rangle}(x))^{\wedge}_{p}$ in Section 2 are fp-spectra of type $n$, the ring spectra $(\operatorname{BP}\langle n\rangle/v_{n}^{e})^{\wedge}_{p}$ is an fp-spectrum of type $n-1$.
@@
Furthermore, we have following theorem establish the relationship between fp-spectra and Lichtenbaum–Quillen property.
Theorem 4.5 ([MR99, Proposition 8.2],[HW22, Theorem 3.1.3]).
@@
is bounded above, where $L^{f}_{n}$ denote the Bousfield localization functor with respect to the spectrum $T(0)\oplus T(1)\oplus\cdots\oplus T(n)$, and $T(k)$ is the telescope of a type $k$ finite complex.
We want to investigate the redshift phenomena in this context, the methods to prove redshift conjecture were well understood, see [AN21, HW22]. More concretely, we have the following theorem. First, we introduce some conventions.
@@
  • $F$-Bounded TR for $X$ = The spectrum $F\otimes\operatorname{TR}(X)$ is bounded.
@@
Note that $\operatorname{TC}\simeq\operatorname{fib}(1-F:\operatorname{TR}\to\operatorname{TR})$, hence we have $F$-Bounded $\operatorname{TR}$Longrightarrow $F$-Bounded $\operatorname{TC}$. In fact, the $\operatorname{TR}$ here is the usual $\operatorname{TR}(-,p)$, and note that we have equivalences $\operatorname{TR}(A^{\wedge}_{p})^{\wedge}_{p}\simeq\operatorname{TR}(A)^{\wedge}_{p}\simeq\operatorname{TR}(A,p)^{\wedge}_{p}$.
@@
$F$-Segal conjecture for $\operatorname{THH}(k[x]/x^{e})^{\wedge}_{p}$Longrightarrow $F$-Bounded $\operatorname{TR}$ for $\operatorname{THH}(k[x]/x^{e})^{\wedge}_{p}$.
@@
First, we recall a framework used in [HW22] to prove the Segal conjecture, and it also appeared in [AK21]. Let $\operatorname{Fil}(\operatorname{Sp}):=\operatorname{Fun}(\mathbb{Z}^{\text{op}},\operatorname{Sp})$ be the category of filtered spectra, which can be equipped with a symmetrical monoidal structure via Day convolution [Lur17, Example 2.2.6.17]. Then for an algebra $X$ in $\operatorname{Fil}(\operatorname{Sp})$, one can define $\operatorname{THH}(x)$ using [NS18, Definition III.2.3.], and $\operatorname{THH}(X)$ is a filtered spectrum. And note that for an algebra $A$, we can easily viewed $A$ as filtered algebra via Whitehead tower which we could denote it by $\tau_{\geq\bullet}A$. Then $\operatorname{THH}(\tau_{\geq\bullet}A)$ gives us a filtration on $\operatorname{THH}(A)$, and the associated spectral sequence is given by
@@
Here, we have used the décalage construction, see [GIKR22, Remark 3.7]. Especially, for $E=\mathbb{F}_{p}$, and $X$ is bounded below, we have the following spectral sequence
@@
Note that according to [Hes18, Proposition 6.6], the Segal conjecture is true for $\operatorname{THH}(\mathbb{F}_{p}[x])$ (and this can be also seen in [HW22]). Hence it would make sense to guess that $F$-Segal conjecture for $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x])$ holds for any type $n+1$-finite complex $F$. And for $\operatorname{BP}\langle n\rangle[x]/x^{e}$, as seen in Proposition 4.9, it would also make sense to guess that Segal conjecture fails for $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e})$. These all require some understanding of the associated graded algebras derived by $\mathbb{F}_{p}$-descent functor ($\mathbb{F}_{p}$-synthetic analogue).
@@
Note that the proof in [Hes18, Proposition 6.6] shows that the Frobenius map $\varphi:\operatorname{THH}(\mathbb{F}_{p}[x])\to\operatorname{THH}(\mathbb{F}_{p}[x])^{tC_{p}}$ is equivalent to the composition of map
@@
which are not in the image of the left hand side, where $j>0,j_{i}\in\{0,1\}$. If we require the total degree greater than $|\sigma x|+\sum_{i=1}^{n}|\sigma a_{i}|=n+1+\sum_{i=1}^{n}|a_{i}|$, such elements does not appear, hence the map modulo $a_{1},a_{2},\ldots,a_{n}$ is an equivalence for $*>|\sigma x|+\sum_{i=1}^{n}|\sigma a_{i}|=n+1+\sum_{i=1}^{n}|a_{i}|$. ∎
This coincides with [Hes18, Proposition 6.6], more generally, one can prove the following theorem for $R=\mathbb{F}_{p}[x_{1},\ldots,x_{m}]$ in Proposition 4.13.
@@
Same as before, and the statement in [Hes18, Proposition 6.6] still works. ∎
@@
is an equivalence in large degrees.
@@
(4.2) $\operatorname{THH}(\operatorname{desc}_{\mathbb{F}_{p}}(\operatorname{BP}\langle n\rangle[x]))^{\wedge}_{v_{0}}\otimes_{\operatorname{desc}_{\mathbb{F}_{p}}\operatorname{BP}\langle n\rangle[x]}\mathbb{F}_{p}[x]\to\operatorname{THH}(\operatorname{desc}_{\mathbb{F}_{p}}\operatorname{BP}\langle n\rangle)^{tC_{p}}\otimes_{\operatorname{desc}_{\mathbb{F}_{p}}(\operatorname{BP}\langle n\rangle[x])}\mathbb{F}_{p}[x]$
@@
$\operatorname{TC}(\operatorname{BP}\langle 1\rangle[x])\otimes\mathbb{S}/p\to L_{2}^{f}(\operatorname{TC}(\operatorname{BP}\langle 1\rangle[x])\otimes\mathbb{S}/p),$
$\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])\otimes\mathbb{S}/p\to L_{2}^{f}(\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])\otimes\mathbb{S}/p).$
@@
${\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])^{\wedge}_{p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$(${\operatorname{TC}(\operatorname{BP}\langle 1\rangle[x])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{K}(\mathbb{Z}_{(p)}[x])^{\wedge}_{p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$((${\operatorname{TC}(\mathbb{Z}_{(p)}[x])}$
Therefore, after applying $F\otimes-$, the above square is still a pull back square. Since $\mathbb{Z}_{(p)}[x]$ is an $\mathbb{E}_{\infty}$-$\operatorname{BP}\langle 0\rangle$-algebra, hence for any $F^{\prime}$ a $p$-local type $2$ finite complex, $F^{\prime}\otimes\operatorname{TC}(\mathbb{Z}_{(p)}[x])$ is bounded. Then $F\otimes\operatorname{TC}(\mathbb{Z}_{(p)}[x])$ is also bounded. Note that K-theory is $\mathbb{A}^{1}$-invariant, then we have $\operatorname{K}(\mathbb{Z}_{(p)}[x])\simeq\operatorname{K}(\mathbb{Z}_{(p)})$. Because $\operatorname{K}(\mathbb{Z}_{(p)})$ is an fp spectrum of type $1$, then $F\otimes\operatorname{K}(\mathbb{Z}_{(p)}[x])$ is bounded. Therefore $F\otimes\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])^{\wedge}_{p}$ is bounded.
The last statement follows from [BHLS23, Lemma 7.22]. ∎
@@
Let $F$ be any $p$-local type $n+2$ finite complex. If $W\not\simeq 0$, then the fiber of the Frobenius
@@
is not bounded above, where $e>1$, $W$ is the fiber of Frobenius map of $F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)$. In other words, $F$-Segal conjecture for $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e})$ is true if and only if $W\simeq 0$. In fact, if $F$ is a sum of generalized Moore complexes $\mathbb{S}/(p^{i_{0}},v_{1}^{i_{1}},\ldots v_{n+1}^{i_{n+1}})$, then $W\not\simeq 0$. Moreover, for $F$ a $p$-local type $n+2$ finite complex, $W\not\simeq 0$.
@@
The unstable Frobenius induces an equivalence
$F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}\otimes B_{m/p}\to F\otimes(\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m})^{tC_{p}}.$
@@
[…]
@@
[…]
@@
[…]
induce isomorphism in large degrees, and the equivalence is due to Lemma 4.21.
@@
Note that $u$ has degree $-2$, weight 00, $|v_{n+1}^{i_{n+1}}|=(2p^{n+1}-2)^{i_{n+1}}$, and it lies in weight $(2p^{n+1}-1)^{i_{n+1}}$. The element $v_{n+1}^{i_{n+1}}$ induce a graded map $\Sigma^{|v_{n+1}^{i_{n+1}}|}\operatorname{gr}(w(v_{n+1}^{i_{n+1}}))\to\operatorname{gr}$. If modulo $v_{n+1}$, $\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p^{i_{0}},\ldots,v_{n}^{i_{n}})$ is connective, then the graded map induce isomorphism on $E_{\infty}$-page at negative degrees. In particular, $E_{\infty}^{-w(v_{n+1}^{i_{n+1}}),-2N-|v_{n+1}^{n+1}|-w(v_{n+1}^{i_{n+1}})}\simeq E_{\infty}^{0,-2N}$, however, when $N\to\infty$, the left hand side is zero (because there is no negative weight elements), while the right hand side is generated by $u^{-N}$. This follows from the following claim.
@@
Let $F$ be any $p$-local type $n+2$ finite complex. If $W\not\simeq 0$, then $F\otimes\operatorname{TR}(\operatorname{BP}\langle n\rangle[x]/x^{e})$ is not bounded, where $W$ is the fiber of the Frobenius map of $F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)$. In fact, $W\not\simeq 0$.

### v3

@@ 3. Negative topological cyclic homology
A spectrum $A$ is height $n$ if $K(h)_{*}A\neq 0$ and $K(h+k)_{*}A=0$ for $k>0$.
@@
Since $\operatorname{BP}\langle n\rangle[x]/x^{e}$ is complex orientated, hence according to [DFH+14, Chapter 6, Proposition 4.1] the localization can be computed as $(v_{m}^{-1}\operatorname{BP}\langle n\rangle[x]/x^{e})^{\wedge}_{(p,v_{1},\ldots,v_{m-1})}$. Therefore for any $m\geq n+1$, $L_{K(m)}\operatorname{BP}\langle n\rangle[x]/x^{e}\simeq 0$ by the fact $v_{m}$ maps to 0 in homotopy, see Remark 2.27. According to Claim 2.9, there is an $\mathbb{E}_{3}$-algebra map $\operatorname{BP}\langle n\rangle[x]/x^{e}\to\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}$. Now by Claim 2.10, there is an $\mathbb{E}_{3}$-algebra map $\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}\to\operatorname{BP}\langle n\rangle$ induced by an $\mathbb{E}_{3}$-$\operatorname{MU}$-algebra map $\operatorname{MU}[x]/x^{e}\to\operatorname{MU}$, using Construction 2.3. Since $\operatorname{K}(-)$ is a lax symmetric monoidal functor [BGT13, Proposition 10.11], and $\operatorname{Mod}_{\operatorname{BP}\langle n\rangle[x]/x^{e}}\to\operatorname{Mod}_{\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}}\to\operatorname{Mod}_{\operatorname{BP}\langle n\rangle}$ are $\mathbb{E}_{2}$-monoidal maps between $\mathbb{E}_{2}$-monoidal categories [LUR17, Proposition 4.5.8.20]. We therefore have following $\mathbb{E}_{2}$-ring maps
@@
With the same notation as in Section 2, $\Lambda_{\operatorname{BP}\langle n\rangle}(x)$ is height $n$, $\operatorname{BP}\langle n\rangle/v_{n}^{e}$ is height n1n-1. And we have
@@
The second $\mathbb{E}_{1}$-$\operatorname{MU}^{hS^{1}}$-algebra map above is induced by an $\mathbb{E}_{1}$-$\operatorname{MU}$-algebra map and applying $^{hS^{1}}$.
@@ 4. Segal conjecture
We investigate Lichtenbaum–Quillen phenomena in this section, we can reduce to proving the Segal conjecture and weak canonical vanishing theorem for certain cyclotomic spectra due to Theorem 4.7. In particular, we prove Proposition 4.16, 4.20. First we review some basic concepts and related techniques.
@@
We say a spectrum $X$ is $\pi$-finite if $X$ is bounded ($\exists a,b\in\mathbb{Z},X\in\operatorname{Sp}_{[a,b]}$), and $\pi_{k}X$ is a finite group for any $k\in\mathbb{Z}$. A spectrum $X$ is called fp-spectrum if $X$ is bounded below, $p$-complete, and there exists a finite complex $F$ such that $F\otimes X$ is finite in the previous sense.
@@
The ring spectra $(\operatorname{BP}\langle n\rangle[x]/x^{e})^{\wedge}_{p},(\Lambda_{\operatorname{BP}\langle n\rangle}(x))^{\wedge}_{p}$ in Section 2 are fp-spectra of type $n$, the ring spectra $(\operatorname{BP}\langle n\rangle/v_{n}^{e})^{\wedge}_{p}$ is an fp-spectrum of type n1n-1.
@@
Furthermore, we have following theorem establishing the relationship between fp-spectra and Lichtenbaum–Quillen property.
Theorem 4.5 ([MR99, Proposition 8.2],[HW22, Theorem 3.1.3][YAN25, Proposition 1.1]).
@@
is bounded above, where $L^{f}_{n}$ denote the Bousfield localization functor with respect to the spectrum $T(0)\oplus T(1)\oplus\cdots\oplus T(n)$, and $T(k)$ is the telescope of a type $k$ finite complex. In fact, if $X$ is a $p$-local spectrum and there is a type-$(n+1)$ finite spectrum $F$ such that $X\otimes F$ is bounded. Then the fiber of the localization map
$X\to L^{f}_{n}X$
is bounded above.
We aim to investigate the redshift phenomena in this context, the methods to prove redshift conjecture were well understood, see [AN21, HW22]. More concretely, we have the following theorem. First, we introduce some conventions.
@@
  • $F$-Bounded TR for XX= The spectrum $F\otimes\operatorname{TR}(X)$ is bounded.
@@
Note that $\operatorname{TC}\simeq\operatorname{fib}(1-F:\operatorname{TR}\to\operatorname{TR})$, hence we have $F$-Bounded $\operatorname{TR}$ $\Longrightarrow$ $F$-Bounded $\operatorname{TC}$. In fact, the $\operatorname{TR}$ here is the usual $\operatorname{TR}(-,p)$, and note that we have equivalences $\operatorname{TR}(A^{\wedge}_{p})^{\wedge}_{p}\simeq\operatorname{TR}(A)^{\wedge}_{p}\simeq\operatorname{TR}(A,p)^{\wedge}_{p}$.
@@
$F$-Segal conjecture for $\operatorname{THH}(k[x]/x^{e})^{\wedge}_{p}$ $\Longrightarrow$ $F$-Bounded $\operatorname{TR}$ for $\operatorname{THH}(k[x]/x^{e})^{\wedge}_{p}$.
@@
First, we recall a framework used in [HW22] to prove the Segal conjecture, and it also appeared in [AS18]. Let $\operatorname{Fil}(\operatorname{Sp}):=\operatorname{Fun}(\mathbb{Z}^{\text{op}},\operatorname{Sp})$ be the category of filtered spectra, which can be equipped with a symmetrical monoidal structure via Day convolution [LUR17, Example 2.2.6.17]. Then for an algebra $X$ in $\operatorname{Fil}(\operatorname{Sp})$, one can define $\operatorname{THH}(x)$ using [NS18, Definition III.2.3.], and $\operatorname{THH}(X)$ is a filtered spectrum. And note that for an algebra $A$, we can easily viewed $A$ as filtered algebra via Whitehead tower which we could denote it by $\tau_{\geq\bullet}A$. Then $\operatorname{THH}(\tau_{\geq\bullet}A)$ gives us a filtration on $\operatorname{THH}(A)$, and the associated spectral sequence is given by
@@
Here, we have used the décalage construction, see [GIK+22, Remark 3.7]. Especially, for $E=\mathbb{F}_{p}$, and $X$ is bounded below, we have the following spectral sequence
@@
Note that according to [HES01, Proposition 6.6], the Segal conjecture is true for $\operatorname{THH}(\mathbb{F}_{p}[x])$ (and this can be also seen in [HW22]). Hence it would make sense to guess that $F$-Segal conjecture for $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x])$ holds for any type $n+1$-finite complex $F$. And for $\operatorname{BP}\langle n\rangle[x]/x^{e}$, as seen in Proposition 4.9, it would also make sense to guess that Segal conjecture fails for $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e})$. These all require some understanding of the associated graded algebras derived by $\mathbb{F}_{p}$-descent functor ($\mathbb{F}_{p}$-synthetic analogue).
@@
Note that the proof in [HES01, Proposition 6.6] shows that the Frobenius map $\varphi:\operatorname{THH}(\mathbb{F}_{p}[x])\to\operatorname{THH}(\mathbb{F}_{p}[x])^{tC_{p}}$ is equivalent to the composition of map
@@
which are not in the image of the left hand side, where $j>0,j_{i}\in\{0,1\}$. If we require the total degree greater than $|\sigma x|+\sum_{i=1}^{n}|\sigma a_{i}|=n+1+\sum_{i=1}^{n}|a_{i}|$, such elements does not appear, hence the map modulo $a_{1},a_{2},\ldots,a_{n}$ is an equivalence for $|\sigma x|+\sum_{i=1}^{n}|\sigma a_{i}|=n+1+\sum_{i=1}^{n}|a_{i}|$. ∎
This coincides with [HES01, Proposition 6.6], more generally, one can prove the following theorem for $R=\mathbb{F}_{p}[x_{1},\ldots,x_{m}]$ in Proposition 4.13.
@@
Same as before, and the statement in [HES01, Proposition 6.6] still works. ∎
@@
If one replaces $\operatorname{BP}\langle n\rangle$ with $\operatorname{BP}\langle n\rangle\otimes\mathbb{S}_{W(k)}$ the statement holds still. Here, $k$ is a perfect field of characteristic $p$, $\mathbb{S}_{W(k)}$ is the spherical Witt vectors.
@@
(4.2) $\operatorname{THH}(\operatorname{desc}_{\mathbb{F}_{p}}(\operatorname{BP}\langle n\rangle[x]))^{\wedge}_{v_{0}}\otimes_{\operatorname{desc}_{\mathbb{F}_{p}}\operatorname{BP}\langle n\rangle[x]}\mathbb{F}_{p}[x]\to\operatorname{THH}(\operatorname{desc}_{\mathbb{F}_{p}}\operatorname{BP}\langle n\rangle[x])^{tC_{p}}\otimes_{\operatorname{desc}_{\mathbb{F}_{p}}(\operatorname{BP}\langle n\rangle[x])}\mathbb{F}_{p}[x]$
@@
$\operatorname{TC}(\operatorname{BP}\langle 1\rangle[x])\to L_{2}^{f}(\operatorname{TC}(\operatorname{BP}\langle 1\rangle[x])),$
$\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])^{\wedge}_{p}\to L_{2}^{f}(\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])^{\wedge}_{p}).$
@@
${\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])^{\wedge}_{p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{TC}(\operatorname{BP}\langle 1\rangle[x])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{K}(\mathbb{Z}_{(p)}[x])^{\wedge}_{p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{TC}(\mathbb{Z}_{(p)}[x])}$
Therefore, after applying $F\otimes$-, the above square is still a pull back square. Since $\mathbb{Z}_{(p)}[x]$ is an $\mathbb{E}_{\infty}$-$\operatorname{BP}\langle 0\rangle$-algebra, hence for any $F^{\prime}$ a $p$-local type $2$ finite complex, $F^{\prime}\otimes\operatorname{TC}(\mathbb{Z}_{(p)}[x])$ is bounded. Then $F\otimes\operatorname{TC}(\mathbb{Z}_{(p)}[x])$ is also bounded. Note that K-theory is $\mathbb{A}^{1}$-invariant for regular rings, then we have $\operatorname{K}(\mathbb{Z}_{(p)}[x])\simeq\operatorname{K}(\mathbb{Z}_{(p)})$. Because $\operatorname{K}(\mathbb{Z}_{(p)})$ is an fp spectrum of type $1$, then $F\otimes\operatorname{K}(\mathbb{Z}_{(p)}[x])$ is bounded. Therefore $F\otimes\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])^{\wedge}_{p}$ is bounded.
The last statement follows from Theorem 4.5. ∎
@@
Let $F$ be any $p$-local type $n+2$ finite complex. Let $W$ be the fiber of Frobenius map of $F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)$. The following statements are true.
  1. (1)
    If $W\not\simeq 0$, then the fiber of the Frobenius
@@
    is not bounded above, where $e>1$. In other words, $F$-Segal conjecture for $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e})$ is true if and only if $W\simeq 0$.
  2. (2)
    If $F$ is a sum of generalized Moore complexes $\mathbb{S}/(p^{i_{0}},v_{1}^{i_{1}},\ldots v_{n+1}^{i_{n+1}})$, then $W\not\simeq 0$.
  3. (3)
    In fact, for $F$ a $p$-local type $n+2$ finite complex, $W\not\simeq 0$.
@@
The Frobenius map of $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e})$ is the product of maps
$\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m/p}\to\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}\otimes B_{m/p}\to(\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m})^{tC_{p}}.$
 [^4] As one can see, one can replace $\operatorname{BP}\langle n\rangle$ by any $\mathbb{E}_{1}$-ring $R$ to get a more general statement for $\operatorname{THH}(R[x]/x^{e})$.
The second map is called unstable Frobenius which is an equivalence
$\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}\otimes B_{m/p}\to\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m})^{tC_{p}}.$
@@
${\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m/p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq\quad}\scriptstyle{\simeq}$${\Phi^{C_{p}}(\operatorname{THH}(\operatorname{BP}\langle n\rangle))\otimes B_{m/p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Phi^{C_{p}}(\operatorname{THH}(\operatorname{BP}\langle n\rangle))\otimes B_{m}^{C_{p}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}\otimes B_{m}^{C_{p}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\Phi^{C_{p}}(\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\operatorname{THH}(\operatorname{BP}\langle n\rangle\otimes B_{m}))^{tC_{p}}}$
@@
${F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}\scriptstyle{\simeq}$${F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e})^{tC_{p}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\prod_{m\geq 0}F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{\prime}}$$\scriptstyle{\operatorname{id}\otimes\varphi\otimes\operatorname{id}}$${F\otimes\prod_{m\geq 0}(\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m})^{tC_{p}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\prod_{m\geq 0}F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}\otimes B_{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota}$${\prod_{m\geq 0}F\otimes(\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m})^{tC_{p}}.}$
@@
${F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{id}\otimes\varphi\otimes\operatorname{id}}$${F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}\otimes B_{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${F\otimes(\operatorname{THH}(\operatorname{BP}\langle n\rangle\otimes B_{mp})^{tC_{p}}}$
induce isomorphism in large degrees, and the equivalence is due to Lemma 4.
@@
Note that $u$ has degree 2-2, weight 0, $|v_{n+1}^{i_{n+1}}|=(2p^{n+1}-2)^{i_{n+1}}$, and it lies in weight $(2p^{n+1}-1)^{i_{n+1}}$. The element $v_{n+1}^{i_{n+1}}$ induce a graded map $\Sigma^{|v_{n+1}^{i_{n+1}}|}\operatorname{gr}(w(v_{n+1}^{i_{n+1}}))\to\operatorname{gr}$. If modulo $v_{n+1}$, $\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p^{i_{0}},\ldots,v_{n}^{i_{n}})$ is connective, then the graded map induce isomorphism on $E_{\infty}$-page at negative degrees. In particular, $E_{\infty}^{-w(v_{n+1}^{i_{n+1}}),-2N-|v_{n+1}^{n+1}|-w(v_{n+1}^{i_{n+1}})}\simeq E_{\infty}^{0,-2N}$, however, when $N\to\infty$, the left hand side is zero (because there is no negative weight elements), while the right hand side is generated by $u^{-N}$. This follows from the following claim.
@@
Let $F$ be any $p$-local type $n+2$ finite complex. Then $F\otimes\operatorname{TR}(\operatorname{BP}\langle n\rangle[x]/x^{e})$ is not bounded.